OpenStudy (anonymous):

Use the formulas a = x^2 – y^2, b = 2xy, c = x^2 + y^2 to answer this question. One leg of a right triangle has length 11 and all sides are whole numbers. Find the lengths of the other two sides. The other leg= And the hypotenuse=

OpenStudy (anonymous):

@dan815 @Preetha @Australopithecus

OpenStudy (nottim):

are each length a particular variable?

OpenStudy (anonymous):

your seeing everything i see... sadly

OpenStudy (er.mohd.amir):

here x2+y2 is must be hypotenuse and other two are legs and x2-y2=11 since if 2xy=11 not satisfy because side are in whole number

OpenStudy (anonymous):

so what is the missing leg and the hypotenuse?

OpenStudy (anonymous):

so 121 is hypotenuse? and what is the other leg?

OpenStudy (er.mohd.amir):

(x2+y2)2=(x2-y2)2+(2xy)2 use Pythagoras theorem put x2-y2=11 gives x4+y4=121+2x2y2

OpenStudy (er.mohd.amir):

find the triplet in which 11 is come like 3,4,5 make right triangle

OpenStudy (anonymous):

i already looked for it, and couldnt find a triplet for the number 11

OpenStudy (nottim):

uh. is it closed now.

OpenStudy (er.mohd.amir):

with 11 a side right triangle is not possible with whole numbers of sides may be Q is misprinted or something else.